Given the vectors 𝐴⃗=5𝑖̂+3𝑗̂+π‘˜Μ‚ and 𝐡⃗⃗=βˆ’3𝑖̂+2π‘—Μ‚βˆ’4π‘˜Μ‚, find the following:

β€’ 𝐴⃗ βˆ™ 𝐡⃗

β€’ Angle between vector A and B

β€’ 𝐴⃗ π‘₯ 𝐡⃗


1
Expert's answer
2022-02-28T10:38:30-0500

(a)

"{\vec A}\cdot {\vec B}=(5𝑖̂+3𝑗̂+𝑘̂)\cdot(−3𝑖̂+2𝑗̂−4𝑘̂)\\
=-15+6-4=-13"

(b)

"\theta=\cos^{-1}\frac{-13}{\sqrt{5^2+3^2+1^2}\sqrt{(-3)^2+2^2+(-4)^2}}\\
=114^\circ"

(c)

"{\vec A}\times {\vec B}=(5𝑖̂+3𝑗̂+𝑘̂)\times(−3𝑖̂+2𝑗̂−4𝑘̂)\\
=10𝑘̂+20𝑗̂+9𝑘̂-12𝑖̂-3𝑗̂-2𝑖̂\\
=-14𝑖̂+17𝑗̂+19𝑘̂"